# Solving Instance Detection from an Open-World Perspective

### v1

@@ Abstract
Instance detection (InsDet) aims to localize specific object instances within a novel scene imagery based on given visual references. Technically, it requires proposal detection to identify all possible object instances, followed by instance-level matching to pinpoint the ones of interest. Its open-world nature supports its wide-ranging applications from robotics to AR/VR, but also presents significant challenges: methods must generalize to unknown testing data distributions because (1) the testing scene imagery is unseen during training, and (2) there are domain gaps between visual references and detected proposals. Existing methods attempt to tackle these challenges by synthesizing diverse training examples or utilizing off-the-shelf foundation models (FMs). However, they only partially capitalize the available open-world information. In this paper, we approach InsDet from an Open-World perspective, introducing our method IDOW. We find that, while pretrained FMs yield high recall in instance detection, they are not specifically optimized for instance-level feature matching. To address this, we adapt pretrained FMs for improved instance-level matching using open-world data. Our approach incorporates metric learning along with novel data augmentations, which sample distractors as negative examples and synthesize novel-view instances to enrich the visual references. Extensive experiments demonstrate that our method significantly outperforms prior works, achieving 10 AP over previous results on two recently released challenging benchmark datasets in both conventional and novel instance detection settings.
@@ 1 Introduction
Instance detection (InsDet) aims to localize object instances of interest in novel scene imagery (Fig. 1), where these objects are specified by some visual references (as known as support templates) [55, 35, 3, 49, 27]. Its open-world nature supports its wide-ranging applications in robotics and AR/VR [25, 64]. For example, a robot can be commanded to search for a customer’s suitcase at an airport [27], or my-keys in my cluttered bedroom. Its open-world nature also presents significant challenges: during training, one has no knowledge of testing data distribution [50], which can be instantiated by the unknown scene imagery [9], novel object instances encountered only in testing [27], and domain gaps between visual references and detected proposals [49].
@@
Status quo. The literature of InsDet has two settings and existing InsDet methods attempt to address the open world in different aspects (Fig. 2). The conventional instance detection (CID) setting provides visual references of object instances for training. Yet, the testing scene is still unknown in training time. Therefore, some methods collect random background images to model diverse testing scenes (Fig. 2a), cut instances from the visual references, and paste them on these background images to construct a synthetic training set [11, 9]. It is worth noting that background image sampling is utilizing open-world information in a way of data collection in the open world. On the other hand, the novel instance detection (NID) setting, introduced recently [27], studies detecting novel object instances that are specified only in testing time where a trained model is not allowed to be tuned further. To approach NID, [27] uses large-scale synthetic data (Fig. 2b) to learn 3D voxel representations from visual references and a matching function between 2D proposals and 3D voxels. Interestingly, [49] uses off-the-shelf pretrained foundation models (Fig. 2c) for InsDet (we term this method OTS-FM). Specifically, OTS-FM uses SAM [22] to detect object proposals in scene imagery, and DINOv2 [39] to represent visual references and proposals for instance-level feature matching. OTS-FM is a non-learned method and hence is applicable in both CID and NID settings. Importantly, it achieves the state-of-the-art performance on the recently released benchmark of HR-InsDet [49].
@@
The open-world perspective. As illustrated in Fig. 1, the challenges of InsDet lie in its open-world nature: the unknown testing data distribution, and domain gaps between visual references and detected proposals. To address the open-world challenges, existing methods take different strategies to exploit the open-world information (Fig. 2a-c). For example, to address the unknown testing scene imagery, some methods sample diverse background images [11, 9], which is a practice of sampling data in the open world. To mitigate the domain gaps between detected proposals and visual references, [27] attempts to learn features for instance-level matching by utilizing external datasets (ShapeNet [5] and ABO [8]). This data utilization can be considered as data collection in the open world. For better proposal detection and more discriminative feature representations, [49] adopts off-the-shelf pretrained foundation models (FMs) SAM [22] and DINOv2 [39], respectively. One can think of FMs as being pretrained in the open world. Different from existing methods which have not fully capitalized the open-world information, we take the open-world perspective as the first principle and solve InsDet by exploiting the Open World (Fig. 2d). We call our approach IDOW.
@@
Insights and novelties. Our insights are rooted in the understanding of InsDet’s open-world nature and the practice of existing methods. Existing methods address open-world issues to some extent (Fig. 2), particularly by sampling data in the open world [9, 27] and adopting foundation models (FMs) pretrained in the open world [49]. These motivate us to develop InsDet methods in the open world (IDOW). We find that open-world detectors yield high recall on object instances (as reported in [49, 27]) but FMs are not tailored to instance-level matching between proposals and visual references. That said, directly using an FM for InsDet is sub-optimal. Therefore, we propose to adapt FMs to instance-level feature representation by metric learning. We propose novel and simple techniques for data augmentation to enhance FM adaptation and feature learning: distractor sampling, and multi-view synthesis. The former helps learn more discriminative features for instance comparison, and the latter enhances visual references particularly when they are few. With these techniques, our IDOW achieves significant boosts over previously reported results on recently released challenging benchmarks (Table 1 and 2).
@@ 2 Related Work
Instance detection. Early methods use local features such as SIFT [32] and SURF [2] from visual references [44] for matching image regions to localize instances [16]. Recent methods train deep neural networks for instance-agnostic proposal detection and proposal-instance matching, achieving significant improvements [35, 27, 49].
@@
In the conventional instance detection (CID) setting, visual references of objects are given during training. To approach CID, prevalent methods synthesize training data in a cut-paste-learn strategy [11, 9]: cutting instances from visual references and pasting them on random background images, then learning a detector on such synthetic data. Other works strive to obtain more training samples by rendering realistic instance examples [19, 17], using data augmentation [9] and synthesizing better training images [26, 9, 12]. Moreover, recent literature introduces the novel instance detection (NID) setting [27], which requires detecting novel instances specified only in testing and does not allow for further model finetuning during testing. [27] utilizes a pretrained open-world detector [20] for proposal detection. Nowadays, the open-world perspective advances proposal detection in a way of learning more generalizable detectors and carrying out open-vocabulary recognition [30]. While a foundational open-vocabulary detector can attach a textual description (e.g., a short phrase) to each proposal, simply using it is insufficient for instance-level matching. [49] realizes this issue although foundational detectors can yield nearly perfect recall. As a result, it uses another FM DINOv2 [39] for proposal-instance matching [49]. Contrary to them, we solve InsDet from the open-world perspective, embracing FMs and data in the open world to adapt FMs towards more discriminative instance-level feature representations.
@@
Foundation models (FMs) are pre-trained on web-scale data, which can be thought of as being sampled in the open world. Different FMs are trained on different types of data and can shine in different downstream tasks related to visual perception, natural language processing, or both. As this work concerns about the visual perception, we briefly review Vision-Language Models (VLMs) and Visual Foundation Models (VFMs). VLMs are pretrained on web-scale image-text paired data [45, 18, 29, 30], demonstrating impressive results in various tasks such as zero-shot recognition [42] and visual grounding [53]. VFMs are pretrained primarily on visual data [7, 4, 65, 14, 39, 57, 22, 60] and can yield impressive visual perception results such as proposal or open-set object detection [61, 30]. In InsDet, the mainstream framework requires proposal detection and visual feature matching. In this framework, we find, as well as reported in [49], open-world detectors achieve nearly perfect recall on the instances of interest, owing to that the open-world pretrained detectors generalize quite well in the open world [61, 30, 47]. In contrast, feature matching is the most rewarding avenue to improve InsDet. Therefore, we focus on improving features for InsDet by adapting VFMs using data sampled in the open world. [49] shows that using an off-the-shelf FM DINOv2 [39] for feature matching greatly enhances InsDet detection; yet, FM adaptation has been still under-explored in InsDet FM adaptation is a well-established research theme in many other areas such as few-shot perception [52, 33] and zero-shot perception [13, 41]. In this work, we introduce simple and novel techniques to adapt FMs to improve InsDet.
@@
The open-world data allows pretraining FMs (as explained above). Coping with such data is often challenging. For example, data between training and testing has distribution shifts or domain gaps [31], as clearly instantiated in InsDet by the unknown testing scene imagery, in addition to the domain gaps between (clean) visual references and (occluded) proposals (Fig. 1). Speaking of distribution shifts, some existing methods sample open-world data, demonstrating a need to develop robust models in the open world [15, 23]. Existing InsDet methods also sample open-world data such as diverse background images and instance examples (Fig. 2) [11, 9]. Even so, while open-world data is abundant, most of it is not valuable in terms of training for specific downstream tasks. To this end, data augmentation aims to synthesize more tailored data to enhance models’ robustness and generalization, including geometry-aware augmentation (e.g., random cropping and flipping) [24, 7, 63], photometric-aware augmentation (e.g., color jittering) [58, 21, 59], and robustness-aware augmentation (e.g., adversarial perturbation) [62]. It is worth noting that Neural Radiance Field (NeRF) [36], a recent rendering technique, can be used to synthesize novel-view images. In this work, we particularly adopt NeRF to synthesize novel-view visual references of object instances and use them to adapt FMs towards more tailored features to InsDet. To the best of our knowledge, using NeRF is under-explored in the literature of InsDet, but we find that doing so remarkably improves InsDet performance.
@@ 3 Instance Detection: Protocols and Methods
We first introduce the protocols of InsDet, including the problem definition, evaluation, and two settings. We then briefly review existing methods and present our techniques to enhance foundation model adaptation to InsDet.
@@ 3.1 Protocols
Problem definition. Instance Detection (InsDet) aims to detect specific object instances from a 2D scene image (Fig. 1), where the objects of interest are defined by a “support set” of some visual references [9, 49, 27]. Previous works capture the visual references from multiple camera views, with QR code being pasted to help estimate camera poses (see row-1 in Fig. 3).
@@
Conventional Instance Detection (CID). In this setting, all object instances of interest are pre-defined during training [9]. That said, one can use their visual references as training data to train detectors. Yet, only visual references are available and testing scene data is still unknown. Hence, InsDet methods in CID often choose to sample diverse background images in the open world [9, 12]. This setting simulates application scenarios in AR/VR, for instance, robots should memorize the items concerned by the customer so as to provide better customized services.
Novel Instance Detection (NID). This setting requires InsDet models to detect novel object instances defined only in testing time and does not allow the models to be finetuned further during testing [27].
In this setting, the de facto practice to train InsDet models turns to external data, which can be either synthetic data or those sampled in the open world. This setting simulates some real-world scenarios, for instance, robots must search for a never-before-seen luggage of a customer at an airport.
Support Ground Truth Cut-Paste-Learn [9] OTS-FM [49] IDOW (ours)
3.2 Methods
We first introduce state-of-the-art methods to provide backgrounds, then present our solution of instance detection from an open-world perspective (IDOW), focusing on foundation model adaptation and data augmentation.
3.2.1 The State of the Art
Cut-Paste-Learn (CPL) [9] was proposed in the CID setting, where object instances of interest are defined during training by their visual references. CPL follows the framework of detection-by-recognition, i.e., detecting instance-agnostic proposals on testing images and recognizing which object instance each proposal belongs to. It uses the given visual references of instances to train both the proposal detector and instance classifier, following a contemporary object detector design. Specifically, it synthesizes training images [9] by cutting the foregrounds from visual references, pasting them on random diverse background images, and learning an ordinary object detector such as Faster RCNN [46].
VoxDet [27] is a recent method that also follows the detection-by-recognition framework. It leverages an open-world detector to obtain instance-agnostic proposals. It learns a reconstruction model to 3D-represent visual references, and a matching model for 3D-to-2D matching to recognize proposals based on the 3D representation. It trains the models by utilizing abundant object instances obtained from external datasets, considered open-world data.
Off-the-shelf foundation models (OTS-FM) [49] leverages a pretrained detector to obtain proposals and another model for matching proposals and visual references to recognize proposals. Interestingly, this method does not introduce extra learning steps or use any publicly available data. Hence, it is a method that can work in both CID and NID settings. Importantly, it achieves the state-of-the-art performance on the recent benchmark HR-InsDet [49], demonstrating that using FMs is already a good starting point for InsDet.
@@ 3.2.2 Foundation Model Adaptation
As an open-world detector yields almost perfect recall [49, 27], we focus on adapting a foundation model to enhance feature representations for InsDet. Although a foundational feature model (e.g., DINOv2 [39]) offers feature representations applicable to various downstream tasks, it is not tailored to InsDet. To improve feature representation, one can finetune this FM over relevant data, such as those sampled in the open world [27]. Intuitively, the finetuned feature is expected to match proposals and visual references better if they are from the same object instances; otherwise, distinguish them. Therefore, we propose to use a metric learning loss for FM adaptation. Moreover, we sample distractors from imagery in the open world. The distractors serve as negatives in metric learning.
We denote an FM as , parameterized by . This FM serves as the feature extractor to produce -dimensional features for any given visual reference image or proposal . We finetune to better serve InsDet such that features of reference images from the same instance should present high similarity, otherwise low similarity. We achieve this by finetuning using a simple metric learning loss . Specifically, we construct triplets of training data , where , and represent an anchor visual reference, positive and negative samples, respectively. The training data contains those sampled or synthesized in the open world [27], as well as visual references available in the CID setting. Negative examples are cross-instance visual examples and distractors (detailed later). Below is the loss:

### v2

@@ Abstract
Instance detection (InsDet) aims to localize specific object instances within a novel scene imagery based on given visual references. Technically, it requires proposal detection to identify all possible object instances, followed by instance-level matching to pinpoint the ones of interest. Its open-world nature supports its broad applications from robotics to AR/VR but also presents significant challenges: methods must generalize to unknown testing data distributions because (1) the testing scene imagery is unseen during training, and (2) there are domain gaps between visual references and detected proposals. Existing methods tackle these challenges by synthesizing diverse training examples or utilizing off-the-shelf foundation models (FMs). However, they only partially capitalize the available open-world information. In contrast, we approach InsDet from an Open-World perspective, introducing our method IDOW. We find that, while pretrained FMs yield high recall in instance detection, they are not specifically optimized for instance-level feature matching. Therefore, we adapt pretrained FMs for improved instance-level matching using open-world data. Our approach incorporates metric learning along with novel data augmentations, which sample distractors as negative examples and synthesize novel-view instances to enrich the visual references. Extensive experiments demonstrate that our method significantly outperforms prior works, achieving 10 AP over previous results on two recently released challenging benchmark datasets in both conventional and novel instance detection settings.
@@ 1 Introduction
Instance detection (InsDet) aims to localize object instances of interest in novel scene imagery (Fig. 1), where these objects are specified by some visual references (as known as support templates) [53, 34, 3, 48, 26]. Its open-world nature supports its wide-ranging applications in robotics and AR/VR. For example, a robot can be commanded to search for a customer’s suitcase at an airport [26], or my-keys in my cluttered bedroom [62]. Its open-world nature also presents significant challenges: during training, one has no knowledge of testing data distribution, which can be instantiated by the unknown scene imagery [9], novel object instances encountered only in testing [26], and domain gaps between visual references and detected proposals [48].
(a) background imagery [11, 9]
(b) object images [26]
(c) Foundation Models [48]
(d) Solving Instance Detection in the Open World (ours)
@@
Status quo. The literature of InsDet has two settings and existing InsDet methods attempt to address the open world in different aspects (Fig. 2). The conventional instance detection (CID) setting provides visual references of object instances for training. Yet, the testing scene is still unknown in training time. Therefore, early method Cut Paste Learn (CPL) collects random background images to model diverse testing scenes (Fig. 2a), cuts instances from the visual references, and pastes them on these background images to construct a synthetic training set [11, 9]. It is worth noting that background image sampling is utilizing open-world information in a way of data collection in the open world. On the other hand, the novel instance detection (NID) setting, introduced recently [26], studies detecting novel object instances that are specified only in testing time where a trained model is not allowed to be tuned further. To approach NID, VoxDet [26] uses large-scale synthetic data (Fig. 2b) to learn 3D voxel representations from visual references and a matching function between 2D proposals and 3D voxels. Recent work OTS-FM [48] uses off-the-shelf pretrained foundation models (Fig. 2c) for InsDet. Specifically, it uses SAM [22] to detect object proposals in scene imagery, and DINOv2 [38] to represent visual references and proposals for instance-level feature matching. OTS-FM is a non-learned method and hence is applicable in both CID and NID settings. Importantly, it achieves the state-of-the-art performance on the recently released benchmark HR-InsDet [48].
@@
The open-world perspective. As illustrated in Fig. 1, the challenges of InsDet lie in its open-world nature: the unknown testing data distribution, and domain gaps between visual references and detected proposals. To address the open-world challenges, existing methods take different strategies to exploit the open-world information (Fig. 2a-c). For example, to address the unknown testing scene imagery, some methods sample diverse background images [11, 9], demonstrating a practice of sampling data in the open world. To mitigate the domain gap between detected proposals and visual references, [26] attempts to learn features for instance-level matching by utilizing external datasets (ShapeNet [5] and ABO [8]). This data utilization can be considered as data collection in the open world. For better proposal detection and more discriminative feature representations, [48] adopts off-the-shelf pretrained foundation models (FMs) SAM [22] and DINOv2 [38], respectively. One can think of FMs as being pretrained in the open world. Different from existing methods which have not fully capitalized the open-world information, we take the open-world perspective as the first principle and solve InsDet by exploiting the Open World (Fig. 2d). We call our approach IDOW.
@@
Insights and novelties. Our insights are rooted in the understanding of InsDet’s open-world nature and the practice of existing methods. Existing methods address open-world issues to some extent (Fig. 2), particularly by sampling data in the open world [9, 23, 26] and adopting foundation models (FMs) pretrained in the open world [48]. These motivate us to develop InsDet methods in the open world (IDOW). We find that open-world detectors yield high recall on object instances (see precision-recall curves in Fig. 9 of the Appendix) but FMs are not tailored to instance-level matching between proposals and visual references – directly using an FM for InsDet is sub-optimal. Therefore, we adapt FMs to instance-level feature representation by metric learning. We propose novel and simple techniques for data augmentation to enhance FM adaptation and feature learning: distractor sampling, and multi-view synthesis. The former helps learn more discriminative features for instance comparison, and the latter enhances visual references particularly when they are few. With these techniques, our IDOW achieves significant boosts over previously reported results on recently released challenging benchmarks (Table 1 and 2).
@@ 2 Related Work
Instance detection. Early methods use local features such as SIFT [31] and SURF [2] to match visual references [43] and image regions to localize instances [16]. Recent methods train deep neural networks for instance-agnostic proposal detection and proposal-instance matching, achieving significant improvements [34, 26, 48]. In the conventional instance detection (CID) setting, visual references of object instances are given during training. To approach CID, prevalent method synthesize training data in a cut-paste-learn (CPL) strategy [11, 9]: cutting instances from visual references and pasting them on random background images, then learning a detector on such synthetic data. Other works strive to obtain more training samples by rendering realistic instance examples [19, 17], using data augmentation [9] and synthesizing better training images [25, 9, 12]. Moreover, recent literature introduces the novel instance detection (NID) setting [26], which requires detecting novel instances specified only in testing and does not allow for further model finetuning during testing. VoxDet [26] utilizes a pretrained open-world detector [20] for proposal detection. Nowadays, the open-world perspective advances proposal detection in a way of learning more generalizable detectors and carrying out open-vocabulary recognition [29]. While a foundational open-vocabulary detector can attach a textual description (e.g., a short phrase) to each proposal, simply using it is insufficient for instance-level matching. [48] realizes this issue although foundational detectors can yield nearly perfect recall. As a result, the recent method OTS-FM uses another FM DINOv2 [38] for proposal-instance matching [48]. Differently, we solve InsDet from the open-world perspective, exploiting open FMs and open data to learn more discriminative instance-level feature representations.
@@
Foundation models (FMs) are pre-trained on web-scale data, which can be thought of as being sampled in the open world. Different FMs are trained on different types of data and can shine in different downstream tasks related to visual perception, natural language processing, or both. As this work concerns about visual perception, we briefly review Vision-Language Models (VLMs) and Visual Foundation Models (VFMs). VLMs are pretrained on web-scale image-text paired data [44, 18, 28, 29], demonstrating impressive results in high-level tasks such as visual grounding and image captioning. VFMs are pretrained primarily on visual data [7, 4, 63, 14, 38, 55, 22, 58] and can yield impressive visual perception results such as proposal or open-set object detection [59, 29]. In InsDet, the mainstream framework requires proposal detection and visual feature matching. In this framework, open-world detectors appear to achieve high recall on the instances of interest, owing to that the open-world pretrained detectors generalize quite well in the open world [59, 29, 46]. In contrast, feature matching still remains a challenge in improving InsDet. Therefore, we focus on improving features for InsDet by adapting VFMs using data sampled in the open world. [48] shows that using an off-the-shelf FM DINOv2 [38] for feature matching greatly enhances InsDet detection; yet, FM adaptation has been still under-explored in InsDet, although it is a well studied in other areas such as few-shot perception [51, 32, 30] and zero-shot perception [13, 40, 41]. Our work introduces simple and novel techniques to adapt FMs to improve InsDet.
Support Ground Truth Cut-Paste-Learn [9] OTS-FM [48] IDOW (ours)
@@
The open-world data is publicly available for pretraining FMs. Coping with such data is also often challenging. For example, data between training and testing has distribution shifts or domain gaps [49, 30], as clearly instantiated in InsDet by the unknown testing scene imagery, in addition to the domain gaps between (clean) visual references and (occluded) proposals (Fig. 1). Speaking of distribution shifts, some existing methods sample open-world data, demonstrating a need to develop robust models in the open world [15, 23]. Existing InsDet methods also sample open-world data such as diverse background images and instance examples (Fig. 2) [11, 9]. Even so, while open-world data is abundant, most of it is not directly useful for training models for specific downstream tasks. In contrast, data augmentation aims to synthesize more tailored data to enhance models’ robustness and generalization, including geometry-aware augmentation (e.g., random cropping and flipping) [24, 7, 61], photometric-aware augmentation (e.g., color jittering) [56, 21, 57], and robustness-aware augmentation (e.g., adversarial perturbation) [60]. It is worth noting that Neural Radiance Field (NeRF) [35], a recent rendering technique, can be used to synthesize novel-view images. In this work, we particularly adopt NeRF to synthesize novel-view visual references of object instances and use them to adapt FMs towards more tailored features to InsDet. To the best of our knowledge, using NeRF is under-explored in the literature of InsDet, but we find that doing so remarkably improves InsDet performance.
@@ 3 Instance Detection: Protocols and Methods
We first introduce the protocols of InsDet, including the problem definition, evaluation, and two settings. We then present our techniques to enhance foundation model adaptation to InsDet.
@@ 3.1 Protocols
Problem definition. Instance Detection (InsDet) aims to detect specific object instances from a 2D scene image (Fig. 1), where the objects of interest are defined by a “support set” of some visual references [9, 48, 26]. Previous works capture the visual references from multiple camera views, with QR code being pasted to help estimate camera poses (see visual references in Fig. 1).
@@
Conventional Instance Detection (CID). In this setting, all object instances of interest are pre-defined during training [9]. That said, one can use their visual references as training data to train detectors. Yet, only visual references are available and testing scene data is still unknown. Hence, InsDet methods in CID often choose to sample diverse background images in the open world [9, 12]. This setting simulates application scenarios in AR/VR. For instance, robots should memorize the items concerned by the customer so as to provide better customized services.
Novel Instance Detection (NID). This setting requires InsDet models to detect novel object instances defined only in testing time and does not allow the models to be finetuned further during testing [26]. In this setting, the de facto practice to train InsDet models turns to external data, which can be either synthetic data or those sampled in the open world. This setting simulates some real-world scenarios, for instance, robots must search for a never-before-seen luggage of a customer at an airport.
3.2 Learning Instance Detector in Open World
We present our solution of instance detection from an open-world perspective (IDOW), focusing on foundation model adaptation and data augmentation.
@@ 3.2.1 Foundation Model Adaptation
As an open-world detector yields quite high recall [48, 26], we focus on adapting a foundation model to enhance feature representations for InsDet. Although a foundational feature model (e.g., DINOv2 [38]) offers feature representations applicable to various downstream tasks, it is not tailored to InsDet. To improve feature representation, one can finetune this FM over relevant data, such as those sampled in the open world [26]. Intuitively, the finetuned feature is expected to match proposals and visual references better if they are from the same object instances; otherwise, distinguish them. Therefore, we propose to use a metric learning loss for FM adaptation. Moreover, we sample distractors from imagery in the open world. The distractors serve as negatives in metric learning.
Support Ground Truth VoxDet [26] OTS-FM [48] IDOW (ours)
We denote an FM as , parameterized by . This FM serves as the feature extractor to produce -dimensional features for any given visual reference image or proposal . We finetune to better serve InsDet such that features of reference images from the same instance should present high similarity, otherwise low similarity. We achieve this by finetuning using a simple metric learning loss . Specifically, we construct triplets of training data , where , and represent an anchor visual reference, positive and negative samples, respectively. The training data contains examples sampled or synthesized in the open world [26], as well as visual references available in the CID setting. Negative examples are cross-instance visual examples and distractors (detailed later). Below is the loss:
@@
Hard example sampling. To construct training batches of triplets, it is important to effectively sample negative examples, easy negative examples do not provide sufficient gradients during training [50]. In this work, we adopt a batch-level hard negative sampling strategy. In each training batch, we first sample a reference image as an anchor . We then sample a random example from the same instance as the positive one . We finally sample a hard negative example from the union of the reference images of all the other instances and the distractors (explained in the next subsection). Mathematically, we find hard negative patch with the objective: , where represents the instance ID of the data , and indicates the total number of visual reference images.
3.2.2 The Proposed Data Augmentation